Alexandra, Ethan, and Sarah spent $1,341 to buy their dad a computer. The three of them used their savings to pay for the computer. Ethan paid $180 more than Alexandra. Alexandra paid $342 less than Sarah. How much did Sarah pay?
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

a+n+s=1341,
n=a+180,
a=s-342
a = 273, n = 453, s = 615

s-342 + s-342+180+ s = 1341
3s-642+180=1341
s=615
Stanborn has 134 instead of 1341

Answer by stanbon(75887)

Alexandra, Ethan, and Sarah spent $1,341 to buy their dad a computer. The three of them used their savings to pay for the computer. Ethan paid $180 more than Alexandra. Alexandra paid $342 less than Sarah. How much did Sarah pay?
--------
Equations:
a + e + s = 1341
e = a+180
a = s-342
-------
Solve for "s"::
Since a = s-342
e = (s-342+180) = s-162
------
Substitute for e and for a and solve for "s":
s-342 + s-162 + s = 134
-------
3s - 504 = 134
3s = 638
s = $212.67 (amt. Sarah paid)
===============
Cheers,
Stan H.
===============